OpenStudy (anonymous):

a large jar and a small jar of coffee weigh 750 grams altogether.If 2 of the large jars and 3 of the small jars of coffee weigh 1750 grams altogether how much does each kind weigh

OpenStudy (ranga):

Let L be the weight of the large jar and S be the weight of the small jar. L + S = 750 2L + 3S = 1750 Solve for L and S.

OpenStudy (anonymous):

Use Substitution. L is weight of large jar. S is weight of small jar. Here are your two equations: L + s = 750 3s + 2L = 1750 Get s by itself. s = 750 - L Substitute 750 - L in for s in the other equation. 3(750 - L) + 2L = 1750 2250 - 3L + 2L = 1750 Now we only have one variable. So solve for l. -1L = -500 L = 500 grams Therefore s = 250 grams (750 - 500) The large jar weighs 500 grams and the small jar weighs 250 grams.